A capacitor consisting of two concentric spheres of radius R1 and R2 = 2.00·R1 has a capacitance of C = 5.00 picoFarads and is charged to a potential difference of 80.0 Volts.
a)Calculate the energy stored in the capacitor.

b)Calculate the charge on this capacitor, when the electrical energy stored is 27.7·10-8Joules. 

 c)If the radius of the outer sphere is increased by a factor of 3.00 while keeping the charge on the plates constant, by what factor does the stored energy change? 
Answer #1

0.) ULCV? 1 + (5*1012F) (Bov)2 U= 16*10-95 Now Used Q = /2cU U= 27.7*10-8J Q = 12*5*10-12 * 27.7*10-8 Q : 1.66 *109 c) CzURE